Four years ago Bob was twice as old as his little sister. Five years from now he will six years older than his sister. How old is each now?
Answer by sunny1992(35)

let,
little sister age = X
Bob age= 2X
current age
sister's age = X+4
Bob age = 2X+4
After 5 years
sister age= X+4+5
Bob age = X+4+5+6 (six year older than sister) OR 2x+4+5
so,
X+4+5+6=2x+4+5
X+15=2X+9
X=6
Sister's current age= 6+4= 10
Bob Current age= 12+4=16
for any query contact at kool_sunny92@hotmail.com